If your IRA compounds continuously for the next 20 years, how much interest will you earn if it pays 3.45% interest, the current balance is $82,000, and you make no contributions or withdrawals?
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

P=Poe^rt
=82000*e^(0.0345*20), the exponential product is 0.69
=82000*e^0.69=$163,484.67
The product of 20 and 3.45 is 69, and the rule of 70/72 means this should be almost exactly doubling, which it is.

Answer by ikleyn(52812)